• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Sirca

    Log entry time 11:16:31 on November 5,2007

    Entry number 212653

    keyword=Report on cryotarget recovery

    JP managed to recover the LHe target.

    Reconstruction of events leading to the target vent:

    Channel 1 of the Loop 1 cryo controller (corresponding to the
    bottom-most temperature sensor in the LHe loop) freaked out.
    The temperature value passed on to the PID failed to update
    i.e. it got stuck at 7K, and the heater power was driven up.
    Ultimately the pressure exceeded about 220 psia (this is
    the last value we saw on the screens) and the target vented.
    (Vent valve limit presently set to about 200 psia.)

    Obviously, the Loop 1 cryo-controller was power-cycled prior to the recovery.